# Freezing-in the Axiverse

### v1

@@ 5.1 Hadronic Axiverse
We first consider the hadronic axiverse (Eq. (31)), where the axions couple to the SM only through gauge interactions. Because there are only three such axion-SM operators, we see immediately that we will have at most $\Delta N_{\rm eff}\sim 3\times 0.027$ (for decoupling temperature above the electroweak scale). We expect that the three axions will have decoupling temperatures scaling as $T_{d}\sim 1/\alpha_{i}^{3}$, so that the QCD axion $a_{\rm QCD}$ decouples at the lowest temperature and the hypercharge axion $a_{1}$ decouples at the highest temperature, at least below the unification scale. This is what we observe in Fig. 1, where we plot the decoupling temperatures of the individual axion states as a function of $f_{a}$. We have made use of the axion-gauge interaction rates computed at finite temperature in Ref. Salvio:2013iaa, and assumed the instantaneous decoupling approximation. The space above the solid black line, labeled $T_{d}>f_{a}$, corresponds to the parameter region where the EFT breaks down. We do not make any claims about temperatures larger than $f_{a}$ because new states may be accessible at these energies that could drastically change our predictions. As a simple example, in the single-axion KSVZ model, the heavy quarks could thermalize with the bath and later decay to axions (see, e.g., Ref. Cheek:2023fht). In principle, such additional states could even appear at scales lower than $f_{a}$, so that our EFT breaks down earlier; however, we assume that our EFT is valid up to $f_{a}$.
@@
In Fig. 8, we show the $\Delta N_{\rm eff}$ predictions in the hadronic axiverse as a function of the reheating temperature $T_{\rm RH}$ and $f_{a}$. Observations of NS cooling Buschmann:2021juv and SN 1987A 1990PhR...198....1R; Springmann:2024mjp impose a lower limit of $f_{a}\geq 3.6\times 10^{8}$ GeV and $f_{a}\geq 10^{9}$ GeV, respectively. In the upper left, we show results with $d=6$ operators turned off, assuming that ${\cal N}\geq 3$ (i.e. the number of axion-SM operators). This does not correspond necessarily to a realistic theory, but is meant to be indicative of the $d=5$ effects. The purple area corresponds to the region where the reheat temperature is too low and dark radiation production is negligible. As one moves in the plot to the upper left, axion production increases. The blue region corresponds to the region in which the QCD axion thermalizes with the SM bath ($T_{\rm RH}>T_{d,{\rm QCD}}$, and the last two green regions correspond to those where the $a_{2}$ and $a_{1}$ axions thermalize, respectively (see Fig. 1 for a direct comparison). CMB-S4 is required to probe any part of this parameter space, whereas CMB-HD is needed to probe down to the QCD axion decoupling line. For $f_{a}<10^{12}$ GeV, these telescopes would be sensitive to cosmological scenarios with parametrically small values of $T_{\rm RH}/f_{a}$. On the other hand, for $f_{a}\gtrsim 10^{12}$ GeV, the QCD axion may contribute to the DM abundance via the misalignment mechanism (44). To show this, in Fig. 8 we indicate the values of $f_{a}$ such that the QCD axion can explain the DM abundance for two choices of the initial misalignment angle $\theta_{i}=0.1,1$.
@@
If the SM unifies into a simple GUT, then the axions coupled to each gauge group are aligned Agrawal:2022lsp, although this is not necessarily realized in semi-simple GUTs, such as Pati-Salam or Trinification. Thus, in simple GUTs only one axion interacts with the SM (Eq. (32)), and its interactions at energies below the GUT scale are dominated by QCD. This is why the decoupling temperature for $a_{\rm GUT}$ is nearly identical to the decoupling temperature for the QCD axion in Fig. 1. We see in the upper left panel of Fig. 9 that only one axion can thermalize, leading to a smaller $\Delta N_{\rm eff}=0.027$, which CMB-HD would be required to observe. In the remaining panels of Fig. 9, we include $d=6$ operators and consider different values of $\cal N$}. We observe a similar pattern as Fig. 8, with sensitivity to both $d=6$ interactions and the value of $\cal N$}.
@@ 5.2.1 Anarchic Axiverse
Anarchic axion-fermion couplings (sec. 3.2.1) lead to the largest possible $\Delta N_{\rm eff}$ contributions, as each coupling $c_{F,nm}$ in Eq. (18) takes on an unsuppressed $\mathcal{O}(1)$ value. We show this in Fig. 3, where we show the decoupling temperature $T_{d}$ of each axion state which is coupled to fermions (the axion-gauge couplings lead to $T_{d}$ as in the hadronic case, and are not shown). As expected, the axions which couple to the top (even in a flavor-violating way) have the lowest $T_{d}$, which scales like $1/y^{2}$, for a Yukawa coupling $y$. There are ten such axion states: the five coupled to the third generation of $Q$, and the five coupled to the third generation of $u$, in total eight of which are flavor-violating. As can be seen in the upper left panel of Fig. 10, if $T_{\rm RH}$ is large enough, all ten axions are thermalized at early times. Such a scenario is excluded by the recent ACT data, even for $f_{a}$ close to the GUT scale $\sim 10^{16}$ GeV, although SPA only excludes parameter space for which additional axions are thermalized. Future CMB surveys will improve the current sensitivity by several orders of magnitude. Due to the large flavor-violating couplings in the anarchic axiverse, beam-dump experiments strongly constrain even light axions. The strongest beam-dump constraint comes from NA62 searches for $K^{+}\to\pi^{+}+{\rm invisible}$ NA62:2024pjp. Ref. Guadagnoli:2025xnt finds $f_{a}/c_{Q,12},f_{a}/c_{d,12}\gtrsim 6\times 10^{11}$ GeV, far stronger than the bound from NS cooling. We shade the region excluded by flavor-violation searches in Fig. 10. Additionally, for $\theta_{i}\sim{\cal O}(1)$, the overclosure bound from QCD axion DM production (44) requires $f_{a}<9\times 10^{11}$ GeV. This choice of $\theta_{i}$, along with the above-mentioned Kaon decay bound with $c_{Q,12},c_{d,12}\sim 1$, rules out most of the $T_{\rm RH}-f_{a}$ parameter space, except for the narrow region bounded by the NA62, SPA, and the $\theta_{i}=1$ lines. As expected, if one allows for a moderate fine-tuning of $\theta_{i}=0.1$, the parameter space to the left of the $\theta_{i}=0.1$ line opens up. Similarly, a smaller $c_{Q,12},c_{d,12}$ would move the flavor bounds towards the left, opening up new parameter space as well. In the additional three panels, we show results including dimension-6 operators for ${\cal N}=10$, 30, and 50. As expected, for sufficiently large $\cal N$}, values of $T_{\rm RH}$ near $f_{a}$ are excluded by existing CMB data.
@@
5.2.2 Froggatt-Neilsen Textured Axiverse
The assumption that the Froggatt-Neilsen mechanism solves the flavor puzzle leads to a suppression of flavor-violating axion couplings relative to the anarchic case, by a factor of $\epsilon$ raised to some power. Unsurprisingly, then, the decoupling temperatures of the flavor-violating axion states increase, so that such a scenario predicts a smaller $\Delta N_{\rm eff}$ than the anarchic case discussed above. We see this behavior in Fig. 4. Explicitly, the two axion states coupled to $t_{R}$ and $c_{R}$ (labeled $a_{u}^{32}$ and $a_{u}^{23}$) have $T_{d}$ larger by a factor $\epsilon^{-2}$ relative to anarchy (note, $\epsilon\approx 0.23$), while those coupled to $t_{L}$ and $c_{L}$ (labeled $a_{Q}^{32}$ and $a_{Q}^{23}$) have $T_{d}$ larger by a factor $\epsilon^{4}$. This can be verified by plugging in the explicit charges chosen in Eq. (38).
Because of this suppression, the $\Delta N_{\rm eff}$ constraints are significantly relaxed in Fig. 11. On the other hand, the NA62 constraint on $K^{+}\to\pi^{+}a$ is only suppressed by a factor $\epsilon^{2}$, so that it is still much stronger than NS cooling. We find that SPA excludes part of the parameter space up to $f_{a}\sim 10^{14}$ GeV even if dimension-6 operators are not active. Future CMB experiments can improve constraints by as much as four orders of magnitude in some parts of the parameter space.
Figure 11: As in Fig. 10, but assuming the axion couplings to the SM fermions arise from a Froggatt-Neilsen texture (Eq. (38)). This weakens the NA62 constraint on flavor violation to $f_{a}\gtrsim 2\times 10^{10}$ GeV.
@@ 5.2.3 MFV Axiverse
We finally investigate a scenario in which the axion couplings are described by MFV. As we saw in Sec. 3.2.3, only six axion states are non-trivially coupled to the SM fermions. By construction, flavor violation is extremely small, and beam-dump constraints are far subdominant to flavor-conserving constraints. For $\mathcal{O}(1)$ coupling, the most stringent limit on $f_{a}$ comes from white dwarf (WD) cooling 2015ApJ...809..141H, requiring $f_{a}\gtrsim 6.1\times 10^{9}$ GeV, which of course becomes weaker if the dimensionless couplings are smaller. We see in Fig. 5 that the four states which are coupled to $Q$ and $u$ have low $T_{d}$ through their interactions with the top quark, while the two states coupled to $d$ and $e$ are suppressed by the smaller $y_{b}$ and $y_{\tau}$, respectively. In the upper left panel of Fig. 12, we show the resulting $\Delta N_{\rm eff}$ for $d=5$ operators only, which would require SO to probe. In the remaining three panels, we show ${\cal N}=5$, 8, and 11 where the $d=6$ contributions get successively larger.
@@ 5.3 Contributions from Dimension-6 Operators
As the dimension-6 operators generically couple every axion to the SM, $N_{\rm eff}$ provides an upper limit on $\cal N$}. We find that this upper limit can be much more stringent than that by requiring the species scale Dvali:2007hz $\Lambda_{s}>f_{a}$, corresponding to ${\cal N}\leq(M_{\rm Pl}/f_{a})^{2}$ (up to a model-dependent $\mathcal{O}(1)$ factor we have ignored). The contribution to axion energy density from dimension-6 operators scales like
@@
Figure 13: (Upper Left) Upper limits on $\cal N$} from the species scale (gray) and from $\Delta N_{\rm eff}$ (red), as a function of $T_{\rm RH}$, for $f_{a}=10^{14}$ GeV. (Upper Right) As in the upper left, but as a function of $f_{a}$ for $T_{\rm RH}=10^{14}$ GeV. (Lower Panel) As in the upper panels, but over the full 2D parameter space. The colors correspond to the maximum $\cal N$} allowed by either the species scale or $\Delta N_{\rm eff}$, as labeled.
@@ 6 Conclusions
We find there are a maximum of 44 independent SM couplings at dimension five, so that only 44 independent axion states couple to the SM—although not all interact strongly enough to thermalize, even for the highest reheating temperatures and for flavor-anarchic couplings to fermions. We characterize the constraints and detection prospects for scenarios where axion-fermion couplings are determined by anarchy, Froggatt-Neilsen textures, minimal flavor violation, and where fermion couplings are absent at tree-level (like in KSVZ-type scenarios). Axion-fermion couplings may be organized by some other principle; to our knowledge no calculation of those couplings has been performed in explicit string constructions as in Gendler:2023kjt for the axion-photon coupling. Such a calculation would be very valuable in narrowing down the expected freeze-in abundance of string theory axions. The dimension-6 operators, on the other hand, are quadratic in the axion fields and can freeze-in the full axion spectrum. There too, no top-down calculation exists, and in our EFT treatment we assume $\mathcal{O}(1)$ Wilson coefficients.
@@
In computing $\Delta N_{\rm eff}$ from the gauge sector and from the dimension six operators, we have used a simplified framework based on the Boltzmann equations for the axion number density that sets the abundance to its thermal value above $T_{d}$, and scales it down by appropriate powers of $\Gamma/H|_{T=T_{\rm RH}}$ below $T_{d}$. This approach leads to $\mathcal{O}(1)$ errors in the vicinity of $T_{d}$. We leave a numerical evaluation of the phase-space density Boltzmann equations (76) and (83) for future work. We additionally ignored thermal corrections to the interaction rates (except in the gauge sector), although we note that our kinetic computations of the interaction rates due to flavor-conserving fermion couplings agrees with those in Ref. Salvio:2013iaa in the context of thermal field theory.
@@ Acknowledgments
We thank A. Prabhu for collaboration during the early stages of this work. We thank D. Dunsky, N. Gendler, S. Gori, S. Heeba, S. Hong, A. Parikh, N. Rodd, B. Safdi, G. Villadoro, and N. Weiner for useful discussions. We thank S. Gori, M. Kongsore, and A. Prabhu for comments on a draft of the manuscript. JTR is supported by the NSF award PHY-2210498 and the BSF grant 2024091. This work was initiated in part at the Aspen Center for Physics, which is supported by National Science Foundation (NSF) grant PHY-2210452. SK is supported in part by the NSF grant PHY-2310572. SK acknowledges the hospitality of the Kavli Institute for Theoretical Physics (KITP) while the work was in progress. This research was supported in part by the NSF grant PHY-2309135 to KITP. The Center for Computational Astrophysics at the Flatiron Institute is supported by the Simons Foundation.

### v2

@@ 5.1 Hadronic Axiverse
We first consider the hadronic axiverse (Eq. (31)), where the axions couple to the SM only through gauge interactions. Because there are only three such axion-SM operators, we see immediately that we will have at most $\Delta N_{\rm eff}\sim 3\times 0.027$ (for decoupling temperature above the electroweak scale). We expect that the three axions will have decoupling temperatures scaling as $T_{d}\sim 1/\alpha_{i}^{3}$, so that the QCD axion $a_{\rm QCD}$ decouples at the lowest temperature and the hypercharge axion $a_{1}$ decouples at the highest temperature, at least below the unification scale. This is what we observe in Fig. 1, where we plot the decoupling temperatures of the individual axion states as a function of $f_{a}$. We have made use of the axion-gauge interaction rates computed at finite temperature in Ref. Salvio et al. (2014), and assumed the instantaneous decoupling approximation. The space above the solid black line, labeled $T_{d}>f_{a}$, corresponds to the parameter region where the EFT breaks down. We do not make any claims about temperatures larger than $f_{a}$ because new states may be accessible at these energies that could drastically change our predictions. As a simple example, in the single-axion KSVZ model, the heavy quarks could thermalize with the bath and later decay to axions (see, e.g., Ref. Cheek et al. (2024)). In principle, such additional states could even appear at scales lower than $f_{a}$, so that our EFT breaks down earlier; however, we assume that our EFT is valid up to $f_{a}$.
@@
In Fig. 8, we show the $\Delta N_{\rm eff}$ predictions in the hadronic axiverse as a function of the reheating temperature $T_{\rm RH}$ and $f_{a}$. Observations of NS cooling Buschmann et al. (2022) and SN 1987A Raffelt (1990); Springmann et al. (2025) impose a lower limit of $f_{a}\geq 3.6\times 10^{8}$ GeV and $f_{a}\geq 10^{9}$ GeV, respectively. In the upper left, we show results with $d=6$ operators turned off, assuming that ${\cal N}\geq 3$ (i.e. the number of axion-SM operators). This does not correspond necessarily to a realistic theory, but is meant to be indicative of the $d=5$ effects. The purple area corresponds to the region where the reheat temperature is too low and dark radiation production is negligible. As one moves in the plot to the upper left, axion production increases. The blue region corresponds to the region in which the QCD axion thermalizes with the SM bath ($T_{\rm RH}>T_{d,{\rm QCD}}$, and the last two green regions correspond to those where the $a_{2}$ and $a_{1}$ axions thermalize, respectively (see Fig. 1 for a direct comparison). CMB-S4 is required to probe any part of this parameter space, whereas CMB-HD is needed to probe down to the QCD axion decoupling line. For $f_{a}<10^{12}$ GeV, these telescopes would be sensitive to cosmological scenarios with parametrically small values of $T_{\rm RH}/f_{a}$. On the other hand, for $f_{a}\gtrsim 10^{12}$ GeV, the QCD axion may contribute to the DM abundance via the misalignment mechanism (44). To show this, in Fig. 8 we indicate the values of $f_{a}$ such that the QCD axion can explain the DM abundance for two choices of the initial misalignment angle $\theta_{i}=0.1,1$.
@@
If the SM unifies into a simple GUT, then the axions coupled to each gauge group are aligned Agrawal et al. (2022), although this is not necessarily realized in semi-simple GUTs, such as Pati-Salam or Trinification. Thus, in simple GUTs only one axion interacts with the SM (Eq. (32)), and its interactions at energies below the GUT scale are dominated by QCD. This is why the decoupling temperature for $a_{\rm GUT}$ is nearly identical to the decoupling temperature for the QCD axion in Fig. 1. We see in the upper left panel of Fig. 9 that only one axion can thermalize, leading to a smaller $\Delta N_{\rm eff}=0.027$, which CMB-HD would be required to observe. In the remaining panels of Fig. 9, we include $d=6$ operators and consider different values of $\cal N$}. We observe a similar pattern as Fig. 8, with sensitivity to both $d=6$ interactions and the value of $\cal N$}.
@@ 5.2.1 Anarchic Axiverse
Anarchic axion-fermion couplings (sec. 3.2.1) lead to the largest possible $\Delta N_{\rm eff}$ contributions, as each coupling $c_{F,nm}$ in Eq. (18) takes on an unsuppressed $\mathcal{O}(1)$ value. We show this in Fig. 3, where we show the decoupling temperature $T_{d}$ of each axion state which is coupled to fermions (the axion-gauge couplings lead to $T_{d}$ as in the hadronic case, and are not shown). As expected, the axions which couple to the top (even in a flavor-violating way) have the lowest $T_{d}$, which scales like $1/y^{2}$, for a Yukawa coupling $y$. There are ten such axion states: the five coupled to the third generation of $Q$, and the five coupled to the third generation of $u$, in total eight of which are flavor-violating. As can be seen in the upper left panel of Fig. 10, if $T_{\rm RH}$ is large enough, all ten axions are thermalized at early times. Such a scenario is excluded by the recent ACT data, even for $f_{a}$ close to the GUT scale $\sim 10^{16}$ GeV, although SPA only excludes parameter space for which additional axions are thermalized. Future CMB surveys will improve the current sensitivity by several orders of magnitude. Due to the large flavor-violating couplings in the anarchic axiverse, beam-dump experiments strongly constrain even light axions. The strongest beam-dump constraint comes from NA62 searches for $K^{+}\to\pi^{+}+{\rm invisible}$ Cortina Gil and others (2025). Ref. Guadagnoli et al. (2025) finds $f_{a}/c_{Q,12},f_{a}/c_{d,12}\gtrsim 6\times 10^{11}$ GeV, far stronger than the bound from NS cooling. We shade the region excluded by flavor-violation searches in Fig. 10. Additionally, for $\theta_{i}\sim{\cal O}(1)$, the overclosure bound from QCD axion DM production (44) requires $f_{a}<9\times 10^{11}$ GeV. This choice of $\theta_{i}$, along with the above-mentioned Kaon decay bound with $c_{Q,12},c_{d,12}\sim 1$, rules out most of the $T_{\rm RH}-f_{a}$ parameter space, except for the narrow region bounded by the NA62, SPA, and the $\theta_{i}=1$ lines. As expected, if one allows for a moderate fine-tuning of $\theta_{i}=0.1$, the parameter space to the left of the $\theta_{i}=0.1$ line opens up. Similarly, a smaller $c_{Q,12},c_{d,12}$ would move the flavor bounds towards the left, opening up new parameter space as well. In the additional three panels, we show results including dimension-6 operators for ${\cal N}=10$, 30, and 50. As expected, for sufficiently large $\cal N$}, values of $T_{\rm RH}$ near $f_{a}$ are excluded by existing CMB data.
@@
5.2.2 Froggatt-Nielsen Textured Axiverse
The assumption that the Froggatt-Nielsen mechanism solves the flavor puzzle leads to a suppression of flavor-violating axion couplings relative to the anarchic case, by a factor of $\epsilon$ raised to some power. Unsurprisingly, then, the decoupling temperatures of the flavor-violating axion states increase, so that such a scenario predicts a smaller $\Delta N_{\rm eff}$ than the anarchic case discussed above. We see this behavior in Fig. 4. Explicitly, the two axion states coupled to $t_{R}$ and $c_{R}$ (labeled $a_{u}^{32}$ and $a_{u}^{23}$) have $T_{d}$ larger by a factor $\epsilon^{-2}$ relative to anarchy (note, $\epsilon\approx 0.23$), while those coupled to $t_{L}$ and $c_{L}$ (labeled $a_{Q}^{32}$ and $a_{Q}^{23}$) have $T_{d}$ larger by a factor $\epsilon^{4}$. This can be verified by plugging in the explicit charges chosen in Eq. (38).
Because of this suppression, the $\Delta N_{\rm eff}$ constraints are significantly relaxed in Fig. 11. On the other hand, the NA62 constraint on $K^{+}\to\pi^{+}a$ is only suppressed by a factor $\epsilon^{2}$, so that it is still much stronger than NS cooling. We find that SPA excludes part of the parameter space up to $f_{a}\sim 10^{12}$ GeV even if dimension-6 operators are not active. Future CMB experiments can improve constraints by as much as four orders of magnitude in some parts of the parameter space.
Figure 11: As in Fig. 10, but assuming the axion couplings to the SM fermions arise from a Froggatt-Nielsen texture (Eq. (38)). This weakens the NA62 constraint on flavor violation to $f_{a}\gtrsim 2\times 10^{10}$ GeV.
@@ 5.2.3 MFV Axiverse
We finally investigate a scenario in which the axion couplings are described by MFV. As we saw in Sec. 3.2.3, only six axion states are non-trivially coupled to the SM fermions. By construction, flavor violation is extremely small, and beam-dump constraints are far subdominant to flavor-conserving constraints. For $\mathcal{O}(1)$ coupling, the most stringent limit on $f_{a}$ comes from white dwarf (WD) cooling Hansen et al. (2015), requiring $f_{a}\gtrsim 6.1\times 10^{9}$ GeV, which of course becomes weaker if the dimensionless couplings are smaller. We see in Fig. 5 that the four states which are coupled to $Q$ and $u$ have low $T_{d}$ through their interactions with the top quark, while the two states coupled to $d$ and $e$ are suppressed by the smaller $y_{b}$ and $y_{\tau}$, respectively. In the upper left panel of Fig. 12, we show the resulting $\Delta N_{\rm eff}$ for $d=5$ operators only, which would require SO to probe. In the remaining three panels, we show ${\cal N}=5$, 8, and 11 where the $d=6$ contributions get successively larger.
@@ 5.3 Contributions from Dimension-6 Operators
As the dimension-6 operators generically couple every axion to the SM, $N_{\rm eff}$ provides an upper limit on $\cal N$}. We find that this upper limit can be much more stringent than that by requiring the species scale Dvali (2010) $\Lambda_{s}>f_{a}$, corresponding to ${\cal N}\leq(M_{\rm Pl}/f_{a})^{2}$ (up to a model-dependent $\mathcal{O}(1)$ factor we have ignored). The contribution to axion energy density from dimension-6 operators scales like
@@
Figure 13: (Upper Left) Upper limits on $\cal N$} from the species scale (gray) and from $\Delta N_{\rm eff}$ (red), as a function of $T_{\rm RH}$, for $f_{a}=10^{14}$ GeV. The dashed black lines indicate the axion mass such that, if all axions have the same mass $m_{a}$ and initial misalignment angle $\theta_{i}=1$, gives the correct DM abundance via misalignment. This assumes there is no misalignment contribution from QCD, which generically overproduces DM at the shown $f_{a}$. For the smallest ${\cal N}=12$ constrained by $N_{\rm eff}$, one requires $m_{a}\approx 2\times 10^{-12}\,(f_{a}/10^{14}\ {\rm GeV})$ eV. (Upper Right) As in the upper left, but as a function of $f_{a}$ for $T_{\rm RH}=10^{14}$ GeV. (Lower Panel) As in the upper panels, but over the full 2D parameter space. The colors correspond to the maximum $\cal N$} allowed by either the species scale or $\Delta N_{\rm eff}$, as labeled.
@@ 6 Conclusions
We find there are a maximum of 44 independent SM couplings at dimension five, so that only 44 independent axion states couple to the SM—although not all interact strongly enough to thermalize, even for the highest reheating temperatures and for flavor-anarchic couplings to fermions. We characterize the constraints and detection prospects for scenarios where axion-fermion couplings are determined by anarchy, Froggatt-Nielsen textures, minimal flavor violation, and where fermion couplings are absent at tree-level (like in KSVZ-type scenarios). Axion-fermion couplings may be organized by some other principle; to our knowledge no calculation of those couplings has been performed in explicit string constructions as in Gendler et al. (2024) for the axion-photon coupling. Such a calculation would be very valuable in narrowing down the expected freeze-in abundance of string theory axions. The dimension-6 operators, on the other hand, are quadratic in the axion fields and can freeze-in the full axion spectrum. There too, no top-down calculation exists, and in our EFT treatment we assume $\mathcal{O}(1)$ Wilson coefficients.
@@
In computing $\Delta N_{\rm eff}$ from the gauge sector and from the dimension six operators, we have used a simplified framework based on the Boltzmann equations for the axion number density that sets the abundance to its thermal value above $T_{d}$, and scales it down by appropriate powers of $\Gamma/H|_{T=T_{\rm RH}}$ below $T_{d}$. This approach leads to $\mathcal{O}(1)$ errors in the vicinity of $T_{d}$. We leave a numerical evaluation of the phase-space density Boltzmann equations (76) and (83) for future work. We additionally ignored thermal corrections to the interaction rates (except in the gauge sector), although we note that our kinetic computations of the interaction rates due to flavor-conserving fermion couplings agrees with those in Ref. Salvio et al. (2014) in the context of thermal field theory.
@@ Acknowledgments
We thank A. Prabhu for collaboration during the early stages of this work. We thank D. Dunsky, N. Gendler, S. Gori, S. Heeba, S. Hong, A. Parikh, N. Rodd, B. Safdi, G. Villadoro, and N. Weiner for useful discussions. We thank S. Gori, M. Kongsore, and A. Prabhu for comments on a draft of the manuscript, and S. Stelzl for comments on Fig. 13. JTR is supported by the NSF award PHY-2210498 and the BSF grant 2024091. This work was initiated in part at the Aspen Center for Physics, which is supported by National Science Foundation (NSF) grant PHY-2210452. SK is supported in part by the NSF grant PHY-2310572. SK acknowledges the hospitality of the Kavli Institute for Theoretical Physics (KITP) while the work was in progress. This research was supported in part by the NSF grant PHY-2309135 to KITP. The Center for Computational Astrophysics at the Flatiron Institute is supported by the Simons Foundation.